OpenStudy (uri):

Stats help. Q.A bag contains balls,numbered from 1 to 11.Two balls are drawn without replacement.What is the probability that both will be of even numbers.

hartnn (hartnn):

how many even number of balls and how many total balls ?

OpenStudy (uri):

2,4,6,8,10 and 11 balls.

hartnn (hartnn):

ok, so # even balls = 5 # total balls =11 2 balls drawn without replacement means 1st a ball is drawn and we do not put that ball back in the bag and another ball is drawn (similar to last Q of 2 cards drawn)

hartnn (hartnn):

so, when we draw 1st ball, what will be the probability that its even numbered ?

hartnn (hartnn):

P(even # ball) = # of even balls / total # of balls =... ?

OpenStudy (uri):

5/11

hartnn (hartnn):

correct! now how many even balls remain ? how many total balls remain ?

OpenStudy (uri):

4/10

OpenStudy (anonymous):

5/11 is the probability of the first even no, and 4/10 will be of second

OpenStudy (uri):

I already said that xD

OpenStudy (anonymous):

P(A U B) = P(A)+P(B)

OpenStudy (uri):

So 5/11 x 4/10 =0.1818

OpenStudy (uri):

@VV-EFF Ayee I don't like that formula.

OpenStudy (anonymous):

hhmmmm! formula choro! logic banao! ;)

hartnn (hartnn):

5/11 x 4/10 =0.1818 is correct ....

OpenStudy (anonymous):

@hartnn careful.....

OpenStudy (uri):

Thankyou @hartnn bhai. :3

OpenStudy (uri):

and @VV-EFF Wasiq bhai.

OpenStudy (anonymous):

its Wasay bhai not wasiq bhai :P

OpenStudy (uri):

Oh sorry *Wasay my bad :3